OpenStudy (anonymous):

PLEASE HELP ME UNDERSTAND Solve for x: 3|x − 3| + 2 = 14 No solutions x = −1, x = 8.3 x = 0, x = 7 x = −1, x = 7

OpenStudy (anonymous):

First step is to get everything but the absolutes alone. So get everything but \(\left| x-3 \right|\) to the other side.

OpenStudy (anonymous):

so subtract 2?

OpenStudy (anonymous):

One more after that, but yes.

OpenStudy (anonymous):

then divide by 3

OpenStudy (anonymous):

Right! So we'll be left with \[\large \left| x-3 \right| = 4\] Now to get the absolutes, you get the negative and positive. So we can split it to look like \[\large x-3=4~~~and~~~x-3=-4\] Now solve them :)

OpenStudy (anonymous):

add 3 in both equations, but one equation has no solution right, since the absolute valuse cant be negative?

OpenStudy (anonymous):

No the x will be negative, the absolute will cancel it out though :)

OpenStudy (anonymous):

ohk?

OpenStudy (anonymous):

so x=1 or x=-1

OpenStudy (anonymous):

Well plug those two in and see if they work.

OpenStudy (anonymous):

they dont

OpenStudy (anonymous):

Both, or just one?

OpenStudy (anonymous):

just one

OpenStudy (anonymous):

poitive 1 doesnt work??

OpenStudy (anonymous):

I'll show you what went wrong. \[\large x-3=4\to \large x=4+3\]

OpenStudy (anonymous):

ohk

OpenStudy (anonymous):

so x=-1 or x=7

OpenStudy (anonymous):

Correct! Good job :)

OpenStudy (anonymous):

thanks! :)

OpenStudy (anonymous):

\[\huge \color{aqua}N\color{fuchsia}o \space \color{lime}P \color{orange}r \color{blue}o \color{maroon}b \color{red}l \color{olive}e \color{purple}m \ddot\smile \]